Assume the framework of the simple asset market model of the nominal exchange rate (NER) where the equilibrium condition is ensured by limitless arbitrage. You are given the following data for the Home (H) and Foreign (F) interest rates, announced to hold at a horizon of one year, respectively it = 5% and it* = 3%.
a) If you also consider that the corresponding forward rate of the Home currency (HC) with respect to the Foreign currency (FC) is Ft = 1 (using the price or academic quotation, as in the textbook), calculate the equilibrium spot exchange rate St of the HC consistent with Covered Interest Parity (CIP). Explain by how much (in % terms) the HC is expected to depreciate or appreciate.
1
Expert's answer
2019-04-24T10:38:17-0400

St = Ft×(1 + it*)/(1 + it) = 1×1.03/1.05 = 0.98, so the HC is expected to depreciate.